Assume that liver cells are cuboidal in shape, with a side length of 23.0 μm.

i) How many liver cells would fit across a pinhead with a diameter of 0.500 mm?

Number of cells = cells

ii) What is the volume of a liver cell?

Volume of cell =    μm3
Answer #1

Answer i):-

Length, (l) of a liver cell side = 23.0 m = 0.023 mm. (1m = 0.001 mm)

Diameter, (d) of a pinhead = 0.500 mm

Number of liver cells (N) which would fit across a pinhead can be determined by dividing diameter of a pinhead by the side length of the liver cell.

N = d / l

N = 0.500 / 0.023 = 21.74 liver cells

---------------------------------------------------------------

Answer ii):-

Volume of a liver cell = (side length of liver cell)3

Volume of a liver cell = (23 m)3 = 12167 m3

----------------------------------------------------------------------------